MILLIMETER-WAVE SPECTROSCOPY OF COLD RB^{85} ATOMS

\underline{JIANING HAN},JAMIL YASIR,PAUL J. TANNER,D. V. L. NORUM and THOMAS F. GAL-LAGHER, Department of Physics, University of Virginia, McCormick Road, Charlottesville, Virginia 22903.

Cold Rb^{85} atoms were prepared by magneto-optical trap. Millimeter-wave has been used to drive nd to (n-2)f (32 ≤ n ≤ 39) one-photon and nd to (n-1)g (31 ≤ n ≤ 36) two-photon transitions. Quantum defects of f and g states of Rb^{85} were calculated. Full analyses will be presented. ^{a}.

aThis work has been supported by Air Force Office of Scientific Research .